OpenStudy (anonymous):

derivative of f(x) = tan^-1(kx) is k/(1+k^2*x^2) if theta is tan^-1(y/x) find dtheta/dy

OpenStudy (anonymous):

\[f(x) = \tan^{-1} kx\] derative is \[\frac{ k }{ 1+x^2k^2 }\] if \[\phi = \tan^{-1} \frac{ y }{ x }\] find \[\frac{ d \phi }{ dy}\]

OpenStudy (anonymous):

tan^-1(y*x^-1) I have now moved up x and it becomes negative. Because Iäm dervating for dy, I know x^-1 can be written as a constant.

OpenStudy (anonymous):

Since its the same formula as the first one I write k = x^-1 or k = 1/x so \[\frac{ k }{ 1+x^2*k^2 }\] can be written as \[\frac{ 1/x }{ 1+((1/x)^2)*ÿ^2}\] Notice as k is replaced by 1/x, x is replaced by y making x^2 = y^2.

OpenStudy (anonymous):

what should be my next step?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

\[\frac{ x }{ x^2+y^2 }\] is the correct answer, I'm at the final step but when I shorten it down it goes wrong..

hartnn (hartnn):

nothing much to simplify just multiply numerator and denominator by x^2 and you will get the required answer

hartnn (hartnn):

(1/x)*x^2 = x = numerator

hartnn (hartnn):

y^2/x^2 * x^2 = y^2 only

OpenStudy (anonymous):

but the correct answer is x /(x^2 + y^2)

OpenStudy (anonymous):

oh Im so stupid, I solved it, or we both did:)

hartnn (hartnn):

yes